what volume of 12.0 mol/L HCL(aq) is required to remove 224g of iron(iii) oxide?
1
Expert's answer
2017-12-15T04:19:48-0500
Solution:
The reaction equation:
Fe 2 O 3 + 6 HCl = 2 FeCl 3 + 3 H 2 O
Amount of Fe 2 O 3 : 224 / 159.69 = 1.4 mol
Amount of HCl needed: 1.4 ∙ 6 = 8.4 mol
Volume of 12M HCl(aq): 8.4 / 12 = 0.7 L
Answer:
0.7 L